What is the sum of


2 raised to 0 ---------------------------------------2 raised to 2008

Grade:9

1 Answers

aditya suman
13 years ago

let s=20+21+--------------------22008

let s1=21+22+---------------22008

s1 is a geometric progression with a=2,r=2 & n=2008

so, s1=a(rn-1)/r-1

s1=2(22008-1)/2-1 = 22009-2 

now, s=20+ s1=1+22009-2=22009-1.